### v2

# Context Matters: Learning Generalizable Rewards via Calibrated Features

Abstract
A key challenge in reward learning from human input is that desired agent behavior often changes based on context. Traditional methods typically treat each new context as a separate task with its own reward function. For example, if a previously ignored stove becomes too hot to be around, the robot must learn a new reward from scratch, even though the underlying preference for prioritizing safety over efficiency remains unchanged. We observe that context influences not the underlying preference itself, but rather the saliency—or importance—of reward features. For instance, stove heat affects the importance of the robot’s proximity, yet the human’s safety preference stays the same. Existing multi-task and meta IRL methods learn context-dependent representations implicitly—without distinguishing between preferences and feature importance—resulting in substantial data requirements. Instead, we propose explicitly modeling context-invariant preferences separately from context-dependent feature saliency, creating modular reward representations that adapt to new contexts. To achieve this, we introduce calibrated features—representations that capture contextual effects on feature saliency—and present specialized paired comparison queries that isolate saliency from preference for efficient learning. Experiments with simulated users show our method significantly improves sample efficiency, requiring 10x fewer preference queries than baselines to achieve equivalent reward accuracy, with up to 15% better performance in low-data regimes (5–10 queries). An in-person user study (N=12) demonstrates that participants can effectively teach their unique personal contextual preferences using our method, enabling more adaptable and personalized reward learning.
1 Introduction
Imagine a robot helping you in the kitchen. When the stove is off, the robot can pass freely above it to reach the counter. When the stove starts heating up, the robot may keep some distance to prevent its end effector from being damaged by the rising heat. When the stove is on high, the robot must stay an even greater distance. This simple scenario highlights a fundamental challenge for robot learning: contexts change, demanding different robot behaviors even when completing the same task.
Traditional approaches to reward learning would treat these as separate tasks with distinct reward functions, even though they stem from the same underlying preference for safety. We argue this fails to capture a key observation: what often changes across contexts is not the human’s preference (e.g., valuing safety over efficiency), but the relative importance – or saliency – of reward features. In our example, the stove’s temperature doesn’t alter the human’s desire for safety, but rather modulates how much the robot’s stove distance matters in achieving that safety. Moreover, these contextual effects on feature saliency are independent of the specific task preference: whether a human generally prioritizes efficiency or safety, a cold stove renders the robot’s stove distance largely irrelevant. This aligns with hypotheses from cognitive science: humans are thought to construct value hierarchically, integrating features whose importance shifts with context o2021hierarchical . Even when high-level preferences remain stable, contextual changes can reshape feature saliency — a relationship we propose to explicitly model.
Prior work in multi-task and meta IRL has focused on learning transferable representations across tasks yamada2022task ; yu2019meta , and may thereby capture some contextual effects. Yet these approaches can at best implicitly learn how context modulates features, often requiring substantial training data. Our key insight is that robots should instead explicitly learn how context affects feature saliency via targeted human feedback, rather than hoping these patterns will emerge with ever-growing amounts of data.
We propose learning these relationships through calibrated features – representations that incorporate contextual effects on task reward features. Instead of learning separate reward functions for each context, our method learns how contextual elements (e.g. stove temperature) should affect task feature values. For example, when the stove is off, the learned calibrated feature should map any distance from the stove to the same value – reflecting that being 2 or 10 steps away makes no practical difference when there’s no heat to avoid. We learn calibrated features as a function of the base feature and the input state. Therefore, the representations capture how the inherent state context affects the feature saliency. By capturing these patterns of contextual influence individually for each feature, our method enables more efficient transfer to new scenarios, even if task preferences change.
We introduce a novel type of human input to learn calibrated features: contextual feature queries – paired comparison queries that explicitly target how context affects feature saliency. A contextual feature query presents a user with two robot states, and asks them to select the context in which the robot better respects a particular feature. In Fig. 1, for the “laptop distance” feature, the user selects the first state with the empty coffee mug over the second state where the full coffee mug is dangerously close to the laptop. For the “stove distance” feature, the user labels the states as equivalent given in both states the stove is off, meaning stove distance is respected no matter the robot’s position. By focusing human feedback on these contextual relationships, we learn how different contexts reshape task features, enabling efficient adaptation to new tasks with minimal supervision.
In summary, our contributions are three-fold. First, we introduce a formal framework for reward learning that explicitly models contextual effects through calibrated features, which can be reused in other tasks, enabling robots to efficiently adapt to new preferences. Second, we propose a method for learning context-calibrated features using targeted contextual feature queries. Finally, we empirically demonstrate that our approach improves sample efficiency and transferability across three complex reward learning tasks in robotic manipulation, outperforming multi-task baselines. A user study validates the ability of real people to respond to contextual feature queries, and uncovers the existence of diversity of contextual preferences, further emphasizing the importance of sample efficiency.
Figure 1: We propose learning representations that capture how context affects the saliency of reward features. (Step 1) The robot asks for contextual queries to learn context-calibrated features. (Step 2) Learned representations can be used to compose various contextually-adaptive downstream rewards.
2 Related Work
Learning from Human Input. Recent advances in robot learning from human input primarily revolve around two paradigms: Inverse Reinforcement Learning (IRL) ng2000algorithms and Behavior Cloning (BC). BC directly learns a policy by mapping expert behavior to actions conditioned on the current state, making it susceptible to distribution shift. In contrast, IRL infers an underlying reward function that explains observed behavior and generalizes to new states. Various forms of human input for IRL have been explored, including demonstrations abbeel2004apprenticeship ; ziebart2008maximum , preferences christiano2017deep ; sadigh2017active , and corrections bajcsy2017learning ; jain2015learning . Beyond reward inference, prior work has also investigated representation learning from human input. Recent methods have aimed to learn transferrable reward representations, either implicitly learning encodings bobu2023sirl or explicitly learning reward features bobu2022inducing , though both without modeling contextual influences. Other recent work has explored using human input to learn safety constraints in the form of control barrier functions to specify unsafe robot behaviors (e.g., limiting a cup’s tilt angle) yang2024enhancing . Our work is complementary to this work. Instead of enforcing hard safety constraints, we provide more granular control by attenuating feature values rather than imposing strict boundaries.
Modeling Context. Limited prior work explicitly models contextual effects on learned behaviors. Related work ghosal2023contextual models contextual effects as a binary switch determining which features are task-relevant. Rather than switching features on and off, we learn representations that account for the effect of context on saliency in a continuous, gradual way. Other work nishi2020fine ; sodhani2021multi structures multi-task rewards as a function of both features and context. These approaches assume explicit access to task context, in the form of metadata sodhani2021multi or discrete values nishi2020fine . In contrast, we don’t assume explicit access to the task context; instead, we learn which state elements compose the relevant context.
Meta and Multi-Task Reward Learning. Multi-task and meta-IRL approaches aim to learn reward functions that transfer across tasks. Early multi-task IRL methods focused on clustering heterogeneous demonstrations and learn distinct rewards babes2011apprenticeship ; dimitrakakis2012bayesian ; choi2012nonparametric , which prevents information sharing across tasks. Following work has explored more complex reward representations with methods for learning shared task representations nishi2020fine ; yamada2022task . Developments in meta-IRL began with applications of MAML finn2017model to tabular settings xu2019learning and known task distributions gleave2018multi , followed by algorithms that introduce latent task variables for policy conditioning yu2019meta ; seyed2019smile . Recent approaches have addressed long-horizon planning chen2023multi and complex environments where maintaining a distribution over tasks is intractable wang2021meta . Prior work has shown that multi-task pre-training followed by fine-tuning can match or outperform meta-reinforcement learning while being more computationally efficient zhao2022effectiveness , therefore we compare our proposed method to a multi-task baseline yamada2022task . It is worth noting that unlike meta and multi-task reward learning methods, our approach does not require multiple training tasks.
3 Approach
3.1 Problem formulation
We are interested in teaching agents reward functions that adapt to context when executing a task. We denote the agent state as $s\in\mathcal{S}$, where $\mathcal{S}=\mathbb{R}^{d}$. We assume the agent has access to $M$ features $\phi(s):\mathbb{R}^{d}\rightarrow[0,1]$. [^1] We normalize base and calibrated features to $[0,1]$ for training stability, but in general they may lie in $\mathbb{R}^{+}$. These features represent useful task aspects (e.g. laptop distance), which can be hand-crafted ziebart2008maximum , learned bobu2022inducing , or even LLM-generated peng2024adaptive . A person’s desired task behavior naturally varies with contextual elements, like stove temperature or cup fullness. Rather than explicitly identifying which state elements constitute context, which is unrealistic in complex environments, we propose a more flexible approach that learns this relationship automatically. We directly model a person’s reward function as $R_{\theta}(\vec{\phi}(s),s):[0,1]^{M}\times\mathbb{R}^{d}\rightarrow\mathbb{R}$, where $\theta$ describes the function approximator parameters. The role of the learned reward $R_{\theta}$ becomes two-fold; it must capture which state elements comprise the relevant context and how these contextual elements should influence the reward. This enables adaptation to new contexts without requiring them to be explicitly labeled.
3.2 Structuring reward with calibrated features
We propose modeling the effects of context using an intermediate representation on top of each feature. We present calibrated features as mappings from a base feature and state to a calibrated feature value, $\phi_{\psi_{i}}^{\prime}(\phi_{i}(s),s):[0,1]\times\mathbb{R}^{d}\rightarrow[0,1]$, where the $i$-th calibrated feature, a function approximator parameterized by $\psi_{i}$, corresponds with the $i$-th base feature. For notational convenience, we use $\phi_{\psi_{i}}^{\prime}(s)$ to denote $\phi_{\psi_{i}}^{\prime}(\phi_{i}(s),s)$, and $\vec{\phi}_{\psi}^{\prime}(s)$ to represent the vector of calibrated features computed separately for each base feature. The role of the $i$-th calibrated feature function is to capture which elements of the state should affect the saliency of the $i$-th base feature, and how. In our example (Fig. 1), the state may consist of robot joint angles, poses of objects (e.g., a cup, laptop, stove, etc.), and properties of those objects (e.g., stove heat or cup fullness). A learned calibrated feature would take as input the state and a base feature, such as the cup angle, and infer which state elements should condition the cup angle feature value (cup fullness) and how (saliency of cup angle is positively correlated with cup fullness). Importantly, we do not learn a function that outputs feature saliency, but one that conditions an input feature on the state context, producing a new, reshaped feature.
We then learn reward on top of calibrated features, $R_{\theta}(\vec{\phi^{\prime}}_{\psi}(s)):[0,1]^{M}\rightarrow\mathbb{R}$, rather than as a function of base features and state. When the reward is linear, parameters $\theta$ maintain their traditional role quantifying the trade-off between features, while calibrated features capture contextual variation within a feature. Shown in Fig. 1, reward parameters may quantify the trade-off between staying away from hot objects and being careful with liquids near technology, while calibrated features reshape how each feature influences the corresponding higher-level concept, based on the current context.
3.3 Learning calibrated features
We leverage paired comparison queries for learning calibrated features, drawing inspiration from prior work on learning features from human input bobu2022inducing . Unlike traditional comparison approaches that ask users to select the state with the greatest reward, we want to solicit input on how changes in state influence the saliency of an individual feature. We present users with contextual feature queries, where they are shown two states and asked to pick in which state a specific feature is better respected, focusing their attention on that feature alone.
Formally, we collect a set of paired comparisons $\mathcal{D}_{\phi_{i}}=\{(s_{1}^{k},s_{2}^{k},y^{k})\}_{k=1}^{K_{i}}$ for each context-calibrated feature $i\in\{1,\ldots,M\}$, where $K_{i}$ is the number of paired comparisons for feature $i$. Paired comparison labels $y^{k}\in\{0,0.5,1\}$ allow the person to indicate whether the feature is more respected in $s_{1}^{k}\ (y^{k}=1)$, in $s_{2}^{k}\ (y^{k}=0)$, or equivalently across states $(y^{k}=0.5)$. Equivalence queries are crucial for identifying which portions of the state should not affect feature saliency, and contexts in which feature saliency should remain static despite changes in input state value. We split the dataset $\mathcal{D}_{\phi_{i}}$ into two subsets which contain all queries labeled with equivalence, $\mathcal{D}_{\phi_{i}}^{\text{equiv}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\phi_{i}}|y=0.5\}$ and all others, $\mathcal{D}_{\phi_{i}}^{\text{pref}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\phi_{i}}|y\neq 0.5\}$.
Following bobu2022inducing we model human feature comparisons using the Bradley-Terry framework bradley1952rank ,
$P_{\psi_{i}}(s_{1}\succ s_{2})=\frac{\exp(\phi^{\prime}_{\psi_{i}}(s_{1}))}{\exp(\phi^{\prime}_{\psi_{i}}(s_{1}))+\exp(\phi^{\prime}_{\psi_{i}}(s_{2}))}\quad,$ (1)
and utilize a cross-entropy loss over a dataset $D$ of paired comparisons,
$\ell(\psi_{i};D)=\sum_{(s_{1},s_{2},y)\in D}\left[-y\log P_{\psi_{i}}(s_{1}\succ s_{2})-(1-y)\log P_{\psi_{i}}(s_{1}\prec s_{2})\right].$ (2)
The final loss equation becomes,
$\mathcal{L}(\psi_{i})=\frac{1}{|D_{\phi_{i}}|}\left(\lambda_{\text{FE}}\cdot\ell(\psi_{i};D_{\phi_{i}}^{\text{equiv}})+\ell(\psi_{i};D_{\phi_{i}}^{\text{pref}})\right),$ (3)
where $\lambda_{\text{FE}}$ is a hyperparameter that scales the contribution of state pairs labeled with equivalence. Since calibrated features are learned individually, users can specify whether each feature depends on context or not, eliminating unnecessary data collection for contextually-unaffected features.
3.4 Learning rewards with calibrated features
In this work, we focus on reward learning from preferences, but our framework could also be adapted to demonstrations finn2016gcl , corrections bajcsy2017learning , or any other reward learning approaches. We ask the human reward preference queries, presenting them with two states and asking them to select which state they prefer. We collect a set of paired comparisons $\mathcal{D}_{\theta}=\{(s_{1}^{n},s_{2}^{n},y^{n})\}_{n=1}^{N}$, where $N$ is the number of reward comparisons. We enable equivalence labels such that $y^{n}\in\{0,0.5,1\}$, indicates if $s_{1}^{n}$ is preferred $(y^{n}=1)$, if $s_{2}^{n}$ is preferred $(y^{n}=0)$ or if the states are viewed as equivalent $(y^{n}=0.5)$. Again, paired comparisons are split into subsets which contain all queries labeled with equivalence, $\mathcal{D}_{\theta}^{\text{equiv}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\theta}|y=0.5\}$ and all others, $\mathcal{D}_{\theta}^{\text{pref}}=\{(s_{1},s_{2},y)\in\mathcal{D}_{\theta}|y\neq 0.5\}$. We model human reward preferences using the same Bradley-Terry framework (Eq. 1) and employ cross entropy loss (Eq. 2), resulting in the final loss function,
$\mathcal{L}(\psi_{i})=\frac{1}{|D_{\theta}|}\left(\lambda_{\text{RE}}\cdot\ell(\theta;D_{\theta}^{\text{equiv}})+\ell(\psi_{i};D_{\theta}^{\text{pref}})\right),$ (4)
where $\lambda_{\text{RE}}$ is a hyperparameter that scales the contribution of equivalent state pairs.
4 Evaluating impact of calibrated features on reward accuracy
We first evaluate our method using simulated human input to objectively measure reward accuracy against ground truth. Our experiments each consist of two stages: (1) representation learning, and (2) reward learning on these representations. We compare rewards learned on calibrated feature representations versus multi-task baseline representations, investigating whether: (1) explicitly capturing contextual effects using calibrated features yields more accurate rewards than multi-task baselines, and (2) learned calibrated features can be effectively composed to generalize to new test rewards.
Environments. We designed three robotic manipulation environments simulated with PyBullet coumans2016pybullet featuring a Jaco robot in a tabletop setting (Fig. 2). Each environment was constructed to reflect everyday scenarios where human preferences may naturally adapt based on varying environmental conditions and object properties. Each environment contains three task-relevant features (described below and in App. B), which enable rich reward composition further complexified by contextual effects. The continuous state space for each environment is a 23-dimensional vector: 3D end-effector (EE) position and rotation matrix (9 values), 3D positions of three fixed objects (laptop, stove, and human), and two environment-specific contextual elements. For experimental evaluation and simulating human responses, we designate these two state elements as the ground truth contextual factors that influence feature saliency (but don’t assume knowledge of the subset at training). The environments facilitate personal contextual relationships, such as the desire to carry heavier objects closer down to the table, or keep sharper objects farther from the human and pointed away.
Weighted Block. Robot carries a block with variable weight. Features are EE xy-planar stove distance and laptop distance, and EE height above table. Contextual elements are stove heat and block weight.
Cup. Robot carries a cup with variable fullness. Features are EE xy-planar stove distance and laptop distance, and EE cup angle. Contextual elements are stove heat and cup fullness.
Utensil. Robot carries a utensil with variable sharpness. Features are EE xy-planar stove distance and human distance, and utensil point at human. Contextual elements are stove heat and utensil sharpness.
Simulating human response. We simulate human responses by defining ground truth calibrated feature functions and reward models. We define seven ground truth calibrated feature functions (App. C) for the seven unique base features across environments, which serve as examples of desired contextual relationships. Ground truth rewards are linear combinations of calibrated features, representing tradeoffs such as it being more important to keep sharp objects away from a person than to avoid heating up the EE. We simulate human responses to state comparison queries using the Bradley-Terry model of human decision making bradley1952rank . We consider states equivalent if their calculated calibrated feature values or rewards fall within an empirically determined threshold (App. C).
Baselines. Following prior work showing multi-task pretraining to be equally or more effective than meta-RL zhao2022effectiveness , we chose a multi-task baseline with variants differing in the number of training tasks. We evaluate our method against JointPref, the most recent applicable, architecture-agnostic, multi-task IRL method yamada2022task where the representations are learned by training $N$ reward heads on a shared representation. This method pressures the shared representation to compactly capture information relevant across all training tasks. We adapt their approach to use preference queries for reward learning, and test implementations with $N=\{10,25,50\}$ (JointPref-N), in addition to $N=1$, representing a single training preference with uniformly weighted features (SinglePref).
4.1 Experimental design
We compare our method to multi-task baselines in two experiments where context affects one or all reward features. In both, we assess accuracy of 10 held-out test rewards learned from preference queries on calibrated features or multi-task representations. We generate 10k states per environment (80/20 train/test split repeated over 6 seeds). We evaluate reward accuracy as the average win rate over pairwise test state comparisons against ground truth. See all experiment details in App. C.
Experiment 1: Single contextually-affected feature. We first aim to evaluate whether explicitly learning contextual effects as in our proposed approach has an advantage over implicit capture by baseline methods. We do this by investigating settings where a single contextual relationship is being learned, in other words only one of the three reward features is affected by context. Since we operate on human-interpretable features, we assume users can identify this contextually-affected feature and we use all contextual queries to train the associated calibrated feature. Letting $i$ denote the index of the calibrated feature, we assume the remaining calibrated feature functions are identity mappings, i.e., $\phi_{\psi_{j}}^{\prime}(s)=\phi_{j}(s)$ for $j\neq i$, and leave them unmodified. In contrast, the baseline method implicitly learns contextual effects through general reward preference queries. Each representation learning method is trained on 100 pre-training queries (rationale in App. C).
Experiment 2: Multiple contextually-affected features. We next investigate if learned calibrated features can be composed to learn accurate downstream rewards. To do this, we use settings where all three reward features vary with context. While we can compare directly with the baseline in this setting, it is important to note that unlike the baseline our method can also compose with new downstream features – regardless of whether they are context-calibrated or not. In this setting our method no longer has an advantage in targeting all pre-training queries to a single calibrated feature. Each representation learning method is trained on 300 pre-training queries (rationale in App. C), which our method splits evenly between the three features.
Figure 2: Calibrated features learned using simulated human input. Calibrated features reshape base feature given changes in affecting contextual element. Visualization details in App. C
Figure 3: Experiment 1 results (single contextually-affected feature). Each row represents the results for reward capturing contextual effects on each of the three features in an environment.
Figure 4: Experiment 2 results (multiple contextually-affected features). Each figure represents the results for capturing contextual effects on all features in an environment.
4.2 Results
Qualitative analysis of learned calibrated features. Fig. 2 visually compares base features to learned calibrated features, shown across changes in the relevant contextual elements. Results show that learned representations successfully capture meaningful feature shaping functions that appropriately calibrate features based on context. In the first row, the representation adapts the base feature of table distance to capture the effect of object weight, i.e. the heavier the object, the closer to the table it should be carried. The concept of safe carrying is encoded by capturing the relative tradeoff of states under different weights – when no object is carried, all reshaped feature values similar, whereas as the object weight increases there is a sharper contrast in calibrated feature values at higher and lower positions. A similar trend appears across all rows: increasing the stove heat, the sharpness of the utensil, and the cup fullness proportionally affects the stove distance, the human distance, and the laptop distance, respectively.
Reward accuracy with single calibrated feature. Fig. 3 illustrates the accuracy of rewards learned by our method compared to multi-task baselines when a single feature in each environment is contextually-affected. Our method consistently outperforms the best-performing baseline across all experiments, with up to 15% improvement in low-data regimes (5-10 pref queries). See details in App. C. The top two performing baselines in most cases are JointPref (10) and SinglePref models. This is likely because as representations are trained only on 100 queries in this experiment, each reward head for JointPref (25) and JointPref (50) only receive 4 and 2 queries each, respectively. This limited data could prevent them from learning a meaningful representation. Notably, when optimizing only a single calibrated feature, the ground truth reward function appears simple enough that SinglePref model performs relatively well but still falls short of our method’s performance.
Reward accuracy with calibrated feature composition. Results (Fig. 4) demonstrate that our approach continues to outperform the best multi-task baseline across all environments, though with a narrower performance gap in this more complex setting. JointPref (10) and JointPref (25) emerge as the strongest baseline methods across most environments, with SinglePref maintaining competitive performance. This aligns with expectations, as the increased number of pre-training queries provides JointPref methods more data per reward head, enabling better learning of meaningful representations.
Figure 5: Subjective user ratings and model ranking improved with increasing training queries.
5 Evaluating human teaching of calibrated features
Given demonstration of our approach’s benefits in simulated human experiments, we ask: can real humans effectively respond to contextual feature queries to teach their personal contextual preferences? Since user contextual preferences lack ground truth definitions, participants subjectively evaluate whether labeling increasing amounts of queries leads to better contextual preference capture.
5.1 User Study
We conducted an in-person, IRB-approved study with 12 participants (7F/5M, ages 22-65, avg=29), each paid \$25 for 70-90 minute sessions. Participants taught their contextual preferences for two Utensil environment features, human distance and stove distance, (with discretized contexts for visualization purposes), with feature order counterbalanced. After an orientation on contextual feature queries and a familiarization round (50 queries for laptop distance from Cup environment), participants completed two experimental rounds (100 queries each). See App. D for study details.
Manipulated variable. For each feature, we trained calibrated features using three different datasets: the first 25 queries, first 50, and all 100. Participants inspected and evaluated four unlabeled models using a graphical user interface: these calibrated feature variants and the uncalibrated base feature.
Evaluation metrics. Participants completed a seven-point Likert scale survey, evaluating how well each model captured their personal contextual preference. The 4-item scale combined items adapted from Hoffman’s goal subscale hoffman2019evaluating (S1: The robot perceives accurately what my preferences are., S2: The robot does not understand how I would like the task to be completed.), prior work in the assessment of adaptive agents nikolaidis2017human (S3: The robot’s learned relationship is responsive to my input.), and a task-specific question (S4: The robot’s learned relationship corresponds well with my expectations for how it should behave in different contexts.). Participants also ranked the presented models based on how well each captured their contextual preferences.
Hypothesis. Increasing the number of human-labeled contextual feature queries used in training improves alignment between the learned calibrated features and human preferences.
5.2 Results
We found partial support for our hypothesis: both Likert ratings and average rankings improved with the number of contextual feature queries used to train the calibrated feature (Fig. 5). A two-way repeated measures ANOVA (feature type × number of training queries) revealed a significant main effect of number of queries, $F(3,33)=19.49,p<0.001,\eta^{2}_{p}=0.427$, but no effect of feature type, $F(1,11)=0.09,p=0.77,\eta^{2}_{p}=0.0005$, and no interaction, $F(3,33)=1.33,p=0.28,\eta^{2}_{p}=0.0333$. Bonferroni-corrected post-hoc tests showed significant differences in Likert ratings between all calibrated features and the base feature $(p<0.05)$, and between 25 and 50 queries $(p<0.05)$. Both average Likert scores and average rankings (1 = best, 4 = worst) improved with the number of training queries, though values leveled off between 50 and 100 queries; see App. D for details.
Overall, calibrated features trained with any amount of contextual feature queries were significantly more aligned with user preferences than the base feature ($p<0.05$). Additionally, features trained with 50% of queries showed significantly better alignment than those trained with 25% ($p<0.05$). While further significance was not observed, likely due to high variance across only 12 participants, the subjective ratings trended positively with training query number. Thus, we find general support that increasing contextual feature queries improves alignment with user contextual preferences.
Finally, post-hoc analysis revealed diverse user preferences, which we grouped into shared high-level categories based on user descriptions. While we observe general trends of contextual preference, there is notable individual variation (Fig. 6). For the human distance feature, 6/12 participants preferred a similar general preference while the remaining 6 described 5 different variations of this preference. For the stove distance feature, 7/12 of participants preferred a similar general preference, while 5/12 held unique variants. These results suggest that while crowdsourcing or warm-starting general preferences may be useful, personalization is crucial to account for individual variations.
Figure 6: Examples of two unique user contextual preferences for how utensil sharpness affects EE-human distance. Users successfully train calibrated features that accurately capture their preferences.
6 Conclusion
Discussion and broader impact. Our work demonstrates that explicit contextual modeling contributes to more flexible and efficient deployment of personalized robots in human environments. By targeting data to teach contextual effects, our method uses human input more effectively, producing representations that better generalize to new preferences without requiring multiple training preferences as demanded by multi-task baselines. The compositionality of our method allows calibrated features to be seamlessly incorporated in new tasks with new features without retraining—a modularity absent in multi-task baselines that lowers the barrier to efficient re-tasking of robots.
Our user study shows that real humans can effectively teach and validate calibrated features using contextual feature queries. This transparency supports iterative refinement and appropriate trust calibration, a critical component for responsible deployment of robots. In contrast, contextual effects remain hidden in the opaque representations of baseline methods. Finally, the diverse contextual preferences revealed in our user study further emphasize the need for sample-efficient learning to achieve preference alignment in widespread deployment.
Limitations. Our method faces several limitations. In continuous spaces, pairwise comparison learning with random sampling can fail to cover important regions of the input space. In Fig. 2 we see this issue result in a failure to capture feature irrelevance in specific contexts, such as when the stove is off or the cup is empty. This problem also affects scenarios where function variation only exists in small portions of the state space; when the cup is full, the calibrated laptop distance collapses to assign low values across most states, whereas the correct representation would assign low values except at the point cloud edge. We don’t observe the same failures in the discrete contexts of the user study (Fig. 6). To address this challenge, methods for improving sample efficiency and data spread could be explored, such as active learning techniques or data augmentation strategies. Future work could explore methods for gathering dense data within single important contexts, potentially drawing inspiration from feature traces bobu2022inducing . Finally, our method requires interpretable base features. Recent work has found ways for LLMs to generate reward features as code automatically from raw state peng2024adaptive , which is a complementary approach that could be used to generate a set of base reward features.
Appendix A Implementation details
A.1 Loss regularization
We implemented the paired comparisons loss described in the main paper with an additional regularization term that penalizes the size of the model logits. For training the reward function $R_{\theta}$, this term regularization is,
$\ell_{\text{reg}}(\theta;\mathcal{D})=\sum_{(s_{1},s_{2},\ell)\in\mathcal{D}}(R_{\theta}(s_{1})^{2}+R_{\theta}(s_{2})^{2}).$ (5)
Similarly, for training calibrated feature function $\phi^{\prime}_{\psi_{i}}$, the analogous regularization term is,
$\ell_{\text{reg}}(\psi_{i};\mathcal{D})=\sum_{(s_{1},s_{2},\ell)\in\mathcal{D}}(\phi^{\prime}_{\psi_{i}}(s_{1})^{2}+\phi^{\prime}_{\psi_{i}}(s_{2})^{2}).$ (6)
With this additional regularization, the implemented loss for training the reward function becomes,
$\mathcal{L}(\theta)=\frac{1}{|D_{\theta}|}\left(\lambda_{\text{RE}}\cdot\ell(\theta;D_{\theta}^{\text{equiv}})+\ell(\theta;D_{\theta}^{\text{pref}})+\lambda_{\text{reg}}\cdot\ell_{\text{reg}}(\theta;D_{\theta})\right).$ (7)
Likewise, for training each calibrated feature, the loss function is structured identically,
$\mathcal{L}(\psi_{i})=\frac{1}{|D_{\psi_{i}}|}\left(\lambda_{\text{FE}}\cdot\ell(\psi_{i};D_{\psi_{i}}^{\text{equiv}})+\ell(\psi_{i};D_{\psi_{i}}^{\text{pref}})+\lambda_{\text{reg}}\cdot\ell_{\text{reg}}(\psi_{i};D_{\psi_{i}})\right).$ (8)
In both cases, $\lambda_{\text{reg}}$ represents the same hyperparameter that weights the regularization term relative to the other loss terms.
A.2 Model architectures and hyperparameters
A.2.1 Selection
Architectures and hyperparameters were optimized separately for each representation learning approach. For each method, we first tuned the hyperparameters for representation learning, then tuned the reward hyperparameters on top of frozen learned representations. Reported hyperparameters were tuned, while all others were left as their PyTorch and optimizer defaults. We conducted all tuning experiments in the Weighted Block environment.
For representation learning methods, in order to maintain sample efficiency, we limited our evaluation to relatively simple MLP architectures, ranging from 1-3 layers and 32-128 units. We selected calibrated feature architecture and hyperparameters using model mean squared error (MSE), and multi-task architecture hyperparameters based on average reward accuracy across the training rewards. Hyperparameter tuning was performed exclusively on the multi-task model with 50 training preferences (JointPref (50)) to leverage the most diverse representation learning scenario, and the resulting configuration was applied consistently across multi-task variants.
For reward learning, we learned a linear combination of calibrated features for our method. For the multi-task baseline, we tested several different architectures, both linear combinations and the same simple MLP architectures used for representations. We gave the baseline access to more expressive reward head architectures than ours to provide the best possible advantage. Architecture and hyperparameters were selected based on reward accuracy across 10 test rewards.
For both representation and reward learning methods, we evaluated whether including or excluding equivalence-labeled queries improved performance. When excluded, equivalent queries were skipped and a new query asked. This strategy only improved performance when learning a linear reward over calibrated features.
A.2.2 Representation learning (pre-training)
Representation learning hyperparameters are summarized in Table 1. The JointPref and SinglePref models used MLPs with three 32-unit layers, outputting to a 7-dimensional latent space. Each calibrated feature network also used three 32-unit layers, with a scalar output passed through a softmax output activation function to produce values in $\mathbb{R}^{+}$. Learned calibrated feature output is normalized using the maximum and minimum logits encountered during training, scaling that the calibrated feature output within the range $[0,1]$. For both approaches, we found that using LeakyReLU activations and Xavier initialization (scaled for LeakyReLU) improved training stability. During representation learning, training task reward heads for all multi-task baselines are modeled as linear combinations of latent representations in order to encourage learned representations to compactly capture meaningful task-relevant variation while remaining amenable to simple downstream reward inference.
Table 1: Representation learning hyperparameters.
Model Type Optim. LR Batch Size Weight Decay $\lambda_{\text{reg}}$ $\lambda_{FE/RE}$ Epochs
Calibrated Features Adam $10^{-3}$ 32 0.01 $10^{-4}$ 10 500
Joint/SinglePref Adam $10^{-3}$ 32 0.01 $10^{-4}$ 10 3000
A.2.3 Reward learning
Reward learning hyperparameters are listed in Table 2. For our method, we learned a linear reward over calibrated features, with frozen representations performing best. Additionally, our method performed the best skipping equivalence queries and generating new non-equivalent queries instead, referenced by the null $\lambda_{\text{FE}}$ value. All multi-task variants used a reward network consisting of a single 32-unit MLP layer (with PyTorch default weight initialization and standard ReLU activations). In low-data settings (5-10 preference queries), multi-task baselines with frozen and unfrozen representations performed similarly, but ultimately unfrozen representations performed best as more data was provided (see Fig. 15&16). Multi-task baseline models with unfrozen representations are the models visualized in the main paper figures.
Table 2: Reward learning hyperparameters.
Model Type Optim. LR Batch Size Weight Decay $\lambda_{\text{reg}}$ $\lambda_{FE/RE}$ Epochs
Calibrated Features Adam $10^{-2}$ 32 0 0 - 200
Joint/SinglePref Adam $10^{-4}$ 64 $10^{-3}$ $10^{-3}$ 1 500
A.3 Compute resources
Each experiment was run on 8 CPU-only compute nodes, each equipped with Intel Xeon Platinum 8260 processors (48 cores per node) and 192 GB of RAM (approximately 4 GB per core). On each node, 24 parallel jobs were run using 2 cores per job. No GPUs were used. Most experiments–such as training a multi-task representation across six random seeds for each query count interval–completed within 5-7 minutes using this setup. For the user study, model training and evaluation were conducted locally on a laptop equipped with an Apple M1 chip and 16GB of memory. Training the four small models on this machine for each query response round of the experiment took approximately one minute.
Appendix B Environment details
We define three distinct environments, each characterized by three features as shown in blue in Fig. 7. The state for each environment is represented as a 23-dimensional continuous vector consisting of end-effector (EE) position (3 values: xyz coordinates) and orientation (9 values: rotation matrix), human position (3 values: xyz coordinates), stove position (3 values: xyz coordinates), laptop position (3 values: xyz coordinates), and 2 environment-specific contextual elements. We generate 10k random states per environment by selecting a random robot EE xyz location above the table within the robot’s reachable workspace, then solving inverse kinematics to obtain joint angles without collisions. The positions of other objects (human, stove, and laptop) remain constant across states. The two contextual elements of a state are randomly sampled with continuous values that range between 0 and 1, however our method can work with contextual elements that follow the state initialization generally in $\mathbb{R}$.
For each environment, all features are normalized to fall within the $[0,1]$ range using min-max normalization across the entire 10k-state dataset. For training stability, the state representation is similarly normalized.
In the Weighted Block environment, the features are z distance to table, xy-planar distance to stove, and xy-planar distance to laptop. These are affected by contextual elements of block weight and stove heat. This environment models scenarios where heavier objects may preferably be carried closer to the table and away from the laptop, while generally maintaining distance from a hot stove.
For the Cup environment, the features are cup angle, xy-planar distance to stove, and xy-planar distance to laptop. These are affected by contextual elements of cup fullness and stove heat. This environment captures situations where fuller cups should be carried more upright and away from the laptop, while generally maintaining distance from a hot stove.
In the Utensil environment, the features are utensil point at person, xy-planar distance to stove, and xy-planar distance to human. These are affected by contextual elements of utensil sharpness and stove heat. This environment represents scenarios where sharper utensils should be carried pointed away from and at a safe distance from the human, while generally maintaining distance from a hot stove.
Figure 7: Three experimental environments with features shown in blue.
The six unique feature definitions are detailed below:
  • Table distance measures the vertical distance between the EE and the table surface:
    $f_{\text{table}}(s)=p^{z}_{\text{EE}}-p^{z}_{\text{table}},$ (9)
    where $p^{z}_{\text{EE}}$ is the z-coordinate of the EE position and $p^{z}_{\text{table}}$ is the z-coordinate of the table surface, which is set as a constant and not part of the state space.
  • Laptop distance measures the planar (XY) distance between the EE and the laptop, capped at a maximum value,
    $f_{\text{laptop}}(s)=\min(\|p^{\text{XY}}_{\text{EE}}-p^{\text{XY}}_{\text{laptop}}\|_{2},0.8),$ (10)
    where $p^{\text{XY}}_{\text{EE}}$ denotes the XY-coordinates of the EE and $p^{\text{XY}}_{\text{laptop}}$ denotes the XY-coordinates of the laptop.
  • Human distance measures the planar (XY) distance between the EE and the human, capped at a maximum value,
    $f_{\text{human}}(s)=\min(\|p^{\text{XY}}_{\text{EE}}-p^{\text{XY}}_{\text{human}}\|_{2},1.0)$ (11)
    where $p^{\text{XY}}_{\text{human}}$ denotes the XY-coordinates of the human.
  • Point at human measures how much a utensil held by the robot points toward the human,
    $f_{\text{point}}(s)=\vec{u}^{W}_{\text{Ex}}{}^{\top}\vec{u}^{W}_{\text{EH}}$ (12)
    where
    $\vec{u}^{,W}_{\text{Ex}}=R^{W}_{\text{EE}}\begin{bmatrix}1,\ 0,\ 0\end{bmatrix}^{\top},\quad\vec{u}^{,W}_{\text{EH}}=\frac{p^{W}_{\text{human}}-p^{W}_{\text{EE}}}{\left\|p^{W}_{\text{human}}-p^{W}_{\text{EE}}\right\|_{2}}.$ (13)
    Here, $\vec{u}^{,W}_{\text{Ex}}$ is the x-axis of the end-effector (i.e., the utensil’s "pointing" direction) expressed in the world frame, obtained via the rotation matrix $R^{W}_{\text{EE}}$ from the end-effector to the world frame. The term $\vec{u}^{W}_{\text{EH}}$ describes the unit vector from the end-effector to the human, also in the world frame. The function equals $1$ when the utensil points directly at the human, and $-1$ when it points directly away.
  • Cup angle The coffee feature measures how upright the cup is held:
    $f_{\text{coffee}}(s)=\left[R^{W}_{WE}\right]_{3,1}$ (14)
    where $\left[R^{W}_{WE}\right]_{3,1}$ denotes the element in the third row and first column of the rotation matrix from the EE frame to the world frame. This represents the z-component of the EE’s x-axis in world coordinates. Since the cup’s upright direction is aligned with the EE’s x-axis, when this x-axis aligns with the world z-axis (vertical), the cup angle value approaches 1, indicating maximum "uprightness".
  • Stove distance The stove feature measures the planar (XY) distance between the EE and the stove, capped at a maximum value:
    $f_{\text{stove}}(s)=\min(\|p^{\text{XY}}_{\text{EE}}-p^{\text{XY}}_{\text{stove}}\|_{2},0.8)$ (15)
    where $p^{\text{XY}}_{\text{stove}}$ denotes the XY-coordinates of the stove.
Discrete contexts for visualization. While all simulated human participant experiments use continuous contexts, visualization requires establishing a set of discrete values for each contextual element (such as a discrete set of utensils varying in sharpness). Each contextual element is discretized with different granularities, all evenly spaced between 0 and 1: stove heat (8 discrete values), block weight (16 discrete values), cup fullness (6 values), and utensil sharpness (4 values). These discrete values are used only for the user study and generating visualization states for qualitative feature visualizations, as detailed in Appendix C.5.
Appendix C Simulated human experiments
C.1 Human simulation
C.1.1 Query response
Ground truth rewards used for simulating human responses are defined as linear combinations of the three environment ground truth calibrated features. In the scenario where only a single feature is contextually affected, two of the three calibrated features behave as identity functions, while the contextually affected feature is transformed according to the ground truth functions described in App. C.1.3. In the scenario where all three features are contextually affected, each is calibrated according to its respective ground truth calibrated feature function.
Ground truth linear reward weights, denoted $\theta^{*}$, are sampled from the set $\{-1,-0.5,0,0.5,1\}$, resulting in a total of 125 possible rewards per environment. A set of 10 test rewards is held out from training multi-task baselines and used solely for evaluation. The SinglePref baseline is trained on a single reward where all calibrated features are equally weighted: $\theta^{*}=[1,1,1]$. Each JointPref ($N$) model includes all rewards used in smaller variants; e.g., JointPref (50) contains the 25 rewards from JointPref (25) plus 25 additional ones. All variants include $\theta^{*}=[1,1,1]$ and $N-1$ additional training rewards.
Human preference selection is simulated using the Bradley-Terry model for pairwise comparisons [6]. Given two states, $s_{1}$ and $s_{2}$ the probability of selecting state $s_{1}$ is
$P_{\theta^{*}}(s_{1}\succ s_{2})=\frac{\exp(\beta\cdot R_{\theta^{*}}(s_{1}))}{\exp(\beta\cdot R_{\theta^{*}}(s_{1}))+\exp(\beta\cdot R_{\theta^{*}}(s_{2}))}\quad,$ (16)
where $\beta=20$ is chosen based on empirical validation. The same setup is used to simulate responses to contextual feature queries, except $\phi^{\prime}_{i}$ is used in place of $R_{\theta^{*}}$ for the $i$-th calibrated feature. The value of $\beta$ is kept constant across all query types. Note that this is the same model of human preference selection used during learning (Eq. 1), although during learning we assume $\beta$ is unknown and fix $\beta=1$.
To simulate equivalence selections, we define a threshold such that two states are marked as equivalent if the absolute difference in their ground truth value (either reward or a specific calibrated feature) falls within that threshold. This threshold was empirically determined, as described in App. C.1.2.
Figure 8: Equivalence threshold experiments for calibrated features.
Figure 9: Equivalence threshold experiments for multi-task baselines.
C.1.2 Equivalence threshold selection
To simulate human judgments of equivalence in paired comparisons, we introduce a threshold $\epsilon$ that defines when two states are considered equivalent. Specifically, this threshold determines whether $|f(s_{1})-f(s_{2})|\leq\epsilon$, in which case the simulated human labels the pair as equivalent, i.e., $f(s_{1})=f(s_{2})$. The function $f$ represents any operation on the state, such as a calibrated feature value or the reward. Using a common $\epsilon$ across both types is reasonable, since the ground truth reward is simply a linear combination (with weights with maximum magnitude of 1), of three calibrated features (each with maximum magnitude of 1).
We select $\epsilon$ by evaluating representation learning under three different candidate thresholds in the Weighted Block environment with all features calibrated. For our method, we plot the MSE between the learned and ground truth calibrated feature values (Fig 8). For the Multipref models, we plot the Reward Accuracy metric evaluated and averaged across all training rewards (Fig. 9).
We observe nearly identical trends across the three tested thresholds, including at key points where the learned representations are used for downstream tasks (i.e., at 100 queries per calibrated feature and 300 queries for each multi-task baseline). Based on this, we select $\epsilon=0.01$, which is the coarsest of the three thresholds and provides the most realistic approximation of human equivalence judgments.
C.1.3 Ground truth calibrated feature function descriptions
Ground truth calibrated feature functions were defined to represent complex three-dimensional relationships governing how context input should shape base features. Intuitively, picking a static context represents taking a slice of through a defined calibrated feature function, demonstrating how the calibrated feature varies with the input base feature under that context.
We define ground truth calibrated features such that they are between 0 and 1 under expected input conditions, when contextual element $c$ and input feature value $\phi$ are in their expected ranges, $c\in[0,1],\phi\in[0,1]$. Importantly, we define ground truth calibrated feature functions to describe how a single contextual element of the state affects a base feature, however our method can handle any subset of the state contextually affecting a base feature, not just a single element. We do so in this work for simplicity of illustration.
We visualize the ground truth calibrated features as a function of the base feature and affecting contextual element in Fig. 10. The actual functions being learned by calibrated feature networks are more complex, given they must not only learn the visualized relationships, but additionally invariance to all other non-relevant dimensions of the state. Descpriptions of the calibrated features for each environment are as follows (with stove distance as affected by stove heat found in all three environments):
  • All environments include the stove distance feature, which is contextually affected by stove heat. The stove distance is evenly respected across all states when the stove is off, however at only the slightest increase in temperature, only a wide berth best respects stove distance.
  • Weighted block environment includes the impact of block weight on both the table distance and the laptop distance. These relationships have similar shapes, though with different implications. For table distance, being far away (larger distance values) violates the feature more when the block is heavy. Once within a certain threshold of the table, any weight equally respects the feature value, and for light blocks, any distance is fine. For laptop distance, being close to directly above the laptop is acceptable for lighter objects, while heavier objects require a wider berth.
  • Cup environment describes the effects of cup fullness on both the cup angle and the laptop distance. Both feature functions are defined very similarly and conservatively—in most scenarios where the cup is at all full, only the widest berth fully respects the feature. The laptop distance is slightly less conservative, demonstrating that for less full cups, there is a slightly larger acceptable berth that fully respects the laptop distance constraint. Only when the cup is empty are both features equally respected everywhere.
  • Utensil environment showcases the result of utensil sharpness on both the point at the human and the distance to the human. As defined, the utensil point calibrated feature deters the utensil from ever being fully pointed at the human, even when dull. There is a sharp cutoff of about halfway, when the utensil points at 90 degrees to the person, when the feature is fully respected. When the utensil is pointing slightly at the person, the feature is more respected when the utensil is dull. The human distance calibrated feature demonstrates a preference for a barrier to be maintained away from the person, regardless of he utensil. This barrier is closer if the utensil is duller, and once past a certain sharpness this barrier is all the way at the edge of the reachable space.
Figure 10: Ground truth calibrated feature functions with respect to base feature and affecting contextual element.
C.1.4 Ground truth calibrated feature function definitions
To help define calibrated feature functions, we first define a handful of base functions that are used to compose the ground truth calibrated feature functions:
  1. 1.
    Gaussian function:
    $G(\phi,c;\sigma,\mu_{\phi},\mu_{c},N,b)=\frac{N}{2\pi\sigma^{2}}\exp\left(-\frac{(\phi-\mu_{\phi})^{2}+(c-\mu_{c})^{2}}{2\sigma^{2}}\right)+b,$ (17)
    which describes a bell-shaped surface with peak at $(\mu_{\phi},\mu_{c})$, decaying exponentially in all directions. Height controlled by $N$, spread by $\sigma$, and vertical shift by $b$.
  2. 2.
    Logistic function:
    $L(\phi;L,k,\mu_{\phi},b)=\frac{L}{1+e^{-k(\phi-\mu_{\phi})}}+b,$ (18)
    which describes an S-shaped curve centered at $\mu_{\phi}$, with steepness $k$, maximum value $L$, and vertical shift $b$.
  3. 3.
    Bowl function:
    $B(\phi,c;N_{\phi},N_{c},\mu_{\phi},\mu_{c},b)=N_{\phi}(\phi-\mu_{\phi})^{2}+N_{c}(c-\mu_{c})^{2}+b,$ (19)
    which generates a parabolic bowl with minimum at $(\mu_{\phi},\mu_{c})$, stretching factors $N_{\phi}$ and $N_{c}$, and vertical shift $b$.
  4. 4.
    Modified logistic:
    $X(\phi;L,k,\mu_{\phi},b)=\frac{L}{1+e^{-k}(\phi-\mu_{\phi})}+b,$ (20)
    which produces a modified logistic function with denominator scaling, creating asymmetric S-curve.
All composed functions follow the general form:
$f_{i}(\phi,c)=\min(\max(f_{\text{base}}(\phi,c),0),1)$
where $\phi$ is the input base feature, $c$ is the relevant contextual state element, and $f_{\text{base}}$ is a single base function or some combination of them. The calibrated features are defined in this form as follows:
  1. 1.
    Calibrated stove feature:
    $f_{\text{stove}}(\phi,c)=\begin{cases}1&\text{if}\ c=0\\ \min(\max(B(\phi,c;1,2,-1,1,-3),0),1)&\text{if}\ c>0\end{cases}$ (21)
  2. 2.
    Calibrated table feature:
    $f_{\text{table}}(\phi,c)=\min(B(\phi,c;1.8,1.5,1,1,0),1)$ (22)
  3. 3.
    Calibrated laptop feature (as affected by block weight):
    $f_{\text{laptop (weight)}}(\phi,c)=\begin{cases}1&\text{if}\ c=0\ \text{or}\ \phi\geq 1\\ \min(\max(B(\phi,c;2.5,2.5,0,1,-1),0),1)&\text{otherwise}\end{cases}$ (23)
  4. 4.
    Calibrated cup angle feature:
    $f_{\text{cup angle}}(\phi,c)=\begin{cases}1&\text{if }c=0\ \text{or}\ \phi\geq 1\\ \min\left(\max\left(\begin{aligned} &X(\phi;-2,-1.1,2,-0.65)\\ &\quad+B(\phi,c;0.5,1.2,-0.2,1,-1)\end{aligned},0\right),1\right)&\text{otherwise}\end{cases}$ (24)
  5. 5.
    Calibrated laptop feature (as affected by cup fullness):
    $f_{\text{laptop (fullness)}}(\phi,c)=\begin{cases}1&\text{if}\ c=0\ \text{or}\ \phi\geq 1\\ \min(\max(B(\phi,c;2,1.5,0,1,-1.5),0),1)&\text{otherwise}\end{cases}$ (25)
  6. 6.
    Calibrated human feature:
    $f_{\text{human}}(\phi,c)=\begin{cases}1&\text{if }\phi\geq 1\\ \min\left(\max\left(\begin{aligned} &G(\phi,c;0.2,1,0,6,-0.55)\\ &\quad+X(\phi;-2,-1.1,2,-0.4)\end{aligned},0\right),1\right)&\text{otherwise}\end{cases}$ (26)
  7. 7.
    Calibrated point feature:
    $f_{\text{point}}(\phi,c)=\begin{cases}1&\text{if }\phi<\frac{1}{3}\\ \min\left(\max\left(\begin{aligned} &X(\phi;0.2,-1,0.5,0)\\ &\quad+G(\phi,c;0.5,0.4,0,2,-0.5)\end{aligned},0\right),1\right)&\text{otherwise}\end{cases}$ (27)
C.2 Training and evaluation
C.2.1 Training
We use a two-stage training process: first learning intermediate representations from training queries, and then training reward functions on top of those representations. For each of six random seeds, we independently split the environment state set of 10k states into 80% training and 20% test states, resulting in different train/test partitions per seed. For each seed, training queries are used for both representation and reward learning, while test queries are used for evaluation only. During representation and reward training, queries are generated by randomly sampling state pairs from the training states.
C.2.2 Metrics
We use two primary metrics to evaluate representations and rewards: calibrated feature accuracy (MSE) and reward accuracy. We evaluate calibrated feature accuracy by computing the mean squared error (MSE) between predicted and ground-truth feature values across the 2,000 test states per seed. Predicted calibrated feature values are normalized between 0 and 1 based on the min and max logits observed during training. Occasionally, especially with untrained models (e.g., 0 training queries), predictions fall outside the logit bounds seen during training (default min and max bounds of 0 and 1), leading to MSE values greater than 1 (see examples in Fig. 12). To evaluate reward functions, we sample 1,000 state pairs from the test set for each seed. We compute the percentage of correctly predicted pairwise preferences according to the ground-truth reward function. This metric is used for both direct evaluation of reward heads trained during representation learning and for downstream reward functions trained on top of frozen or fine-tuned representations. If the ground-truth rewards of a test comparison state pair are within the selected equivalence threshold (App. C.1.2), that pair is excluded from evaluation. Across all 720 test comparison state sets (comprising 10 test rewards evaluated for the full calibrated features reward and each of the 3 single-feature calibrated rewards, across 3 environments and 6 seeds), the number of non-equivalent comparison pairs used for evaluation (out of 1000 possible) had a median of 1000, mean of 956 (SD = 108), and a minimum of 394. Only about 5% of test sets had fewer than 700 usable comparison pairs, with over 80% of those cases stemming from the single test reward in which only one feature was relevant.
C.2.3 Evaluation
Calibrated feature representation learning. For representations that output calibrated features, we evaluate the learned features using the MSE metric described above. This allows us to assess how accurately each feature dimension captures its intended ground-truth signal, even with varying amounts of training data.
Multi-Task representation learning. For multi-task baselines, we evaluate reward accuracy for each of the $N$ training rewards on the same 1,000 test state pairs (excluding equivalent pairs under the training reward). We then average the reward accuracy across the $N$ heads for each seed. This measures how well the shared representation supports multiple reward functions simultaneously.
Downstream reward learning. To assess generalization, we train 10 new (held-out) reward functions on top of the learned representations. These rewards are trained either with the learned representation frozen or unfrozen. Each reward is evaluated on the same 1,000 test state pairs (excluding equivalent pairs), and we report the average reward accuracy across the 10 test rewards for each seed.
C.2.4 Aggregation and error bars
All results are averaged over six seeds. For reward accuracy, we first average across the evaluated rewards ($N$ multi-task training rewards or 10 downstream test rewards) within each seed, then report the mean and standard error across seeds. Therefore, error bars shown in plots reflect standard error across the seed variation, not reward variation.
C.3 Representation learning results
We first present representation learning results for both calibrated feature models and multi-task baselines. For calibrated features, we plot the mean squared error (MSE) of all three learned calibrated features for each environment as a function of the number of training queries (Fig. 11). Each curve represents the MSE for one calibrated feature, evaluated on held-out test states. We also include a zoomed-out version of these plots (Fig. 12), where a dashed horizontal line indicates the MSE of a randomly initialized model (computed as the average MSE across the three features at zero training queries). Calibrated feature MSE can struggle to approach zero when learning from comparisons, which only provide ordinal relationships between pairs of points rather than absolute function values. Some smoother functions (such as cup and laptop distance, which have similar defined calibrated feature functions) are easier to capture accurately with only pairwise comparisons, while others (e.g., distance to human) can struggle to achieve low MSE (Fig. 11).
We next present multi-task representation learning results, which report reward accuracy averaged over the $N$ training rewards in each environment (Fig. 1314). Notably, for multi-task baselines, separate representation learning was required for each downstream reward combination – for each single calibrated feature scenario and for the all calibrated features scenario – since the shared representation depends on the specific reward training setup. In contrast, our calibrated feature approach supports reuse: the same three calibrated features can be applied across all downstream reward configurations in an environment, enabling flexible generalization without retraining the representation. We can see that as expected, the SinglePref model achieves high reward accuracy the most rapidly, given it is only trained and evaluated on a single reward. In contrast, as the number of $N$ training tasks for each JointPref method increases, the more queries it takes to find a shared representation that supports accurate rewards across all training tasks.
We select 100 pre-training queries given we see general convergence behavior for calibrated features at this point. When all three features are contextually-affected we use 300 pre-training queries so that all calibrated features still see 100 pre-training queries when divided among the three features.
Figure 11: Calibrated feature learning results.
Figure 12: Fig. 11 results compared to random network initialization.
Figure 13: Multi-task representation learning results for multiple contextually-affected feature experiments.
Figure 14: Multi-task representation learning results for single contextually-affected feature experiments.
C.4 Reward learning with and without frozen representations
We visualize downstream reward learning results (Fig. 1516), displaying performance when intermediate representations are either frozen or fine-tuned during reward training. For our method, we observe that freezing the learned calibrated feature representations consistently yields higher reward accuracy, particularly in low-data regimes. This suggests that our method successfully learns meaningful and reusable intermediate features that generalize well to new tasks. In contrast, for the multi-task baseline, freezing and unfreezing perform similarly at low query counts, with frozen representations sometimes performing slightly better. However, as more data becomes available, fine-tuning the representation eventually leads to higher performance, indicating that the multi-task representation benefits from task-specific adaptation. In the main text we show the frozen variant of our method and the unfrozen variants of all multi-task baseline methods.
To quantify the overall performance gap, we compute the difference between the best-performing version of our method (always frozen) and the best-performing multi-task baseline configuration (which varies by $N$ training rewards and freezing condition). We summarize these comparisons in Table 3, reporting the also the relative increase in reward accuracy (increase in percentage of correct query selection) demonstrated by our method at both 5 and 10 reward queries. In these low-data regimes, our method achieves improvements of up to 14.91% in reward accuracy over the highest-performing baseline.
Env. Calib. Feats CF 5q JointPref 5q $\Delta\%$ 5q CF 10q JointPref 10q $\Delta\%$ 10q
Weighted All 0.808 0.742 +6.6 0.848 0.773 +7.5
Block Stove 0.802 0.653 +14.9 0.841 0.698 +14.3
Table 0.804 0.663 +14.1 0.844 0.697 +14.7
Laptop 0.794 0.665 +12.9 0.828 0.705 +12.3
Cup All 0.807 0.714 +9.3 0.824 0.726 +9.8
Stove 0.747 0.653 +9.4 0.813 0.692 +12.1
Cup Angle 0.763 0.660 +10.3 0.835 0.688 +14.7
Laptop 0.814 0.690 +12.4 0.840 0.731 +10.9
Utensil All 0.793 0.754 +3.9 0.823 0.775 +4.8
Stove 0.855 0.733 +12.2 0.867 0.783 +8.4
Human 0.827 0.691 +13.6 0.853 0.762 +9.1
Point 0.802 0.743 +5.9 0.825 0.772 +5.3
Table 3: Reward accuracy comparison between calibrated features and the best-performing multi-task variant at 5 and 10 training queries. $\Delta\%$ indicates percent improvement in reward accuracy.
C.5 Qualitative visualization details
The qualitative visualizations display learned calibrated feature values at 8,000 different EE locations across four distinct contexts. The visualization is limited to 8,000 points due to PyBullet’s rendering constraints. Visualization states are generated similar to how described before in App. B, by sampling random points within the robot’s reachable workspace and solving inverse kinematics to obtain joint angles (with object positions remaining constant). However, rather than using random contexts, each point is evaluated across the discrete contexts defined at the end of App. B. For each visualized feature, the relevant contextual element is systematically varied across its defined discrete values while the calibrated feature is evaluated for all points in the set. Non-relevant contextual elements are held constant at zero.
Feature values are depicted as points at each EE location, where lighter colors indicate higher feature values. Learned calibrated feature values are normalized across the entire point cloud evaluated over all discretized contextual values, with non-affecting contextual features held constant. We then select four spread out contexts to visualize the point cloud evaluation in our final figures. All visualizations use calibrated features trained with seed=0 for consistency. We visualize just four features given these visualizations work well for 3D distance-based relationships, but become more challenging for complex transformations like rotations, which would benefit from projection onto relevant 2D planes.
Figure 15: Multiple contextually-affected feature experiments shown with frozen and unfrozen representations.
Figure 16: Single contextually-affected feature experiments shown with frozen and unfrozen representations.
Appendix D User Study
D.1 Result details
The visualized box and whisker plots in 5 display the distribution of data through five key summary statistics. The box represents the interquartile range (IQR) with the bottom edge at the 25th percentile and top edge at the 75th percentile, while a line inside the box marks the median (50th percentile). The whiskers extend from the box edges to the most extreme data points that fall within 1.5 times the IQR from the box boundaries. Any data points beyond this range are plotted as individual outlier points, shown as small circles.
Further user study statistics can be visualized in Table 4. Statistics are shown for both test features individually, as well as across both features. Items on the 7-point Likert scale were summed to represent a composite score representing general contextual preference alignment, making the possible Likert scale score range 4 to 28, given 4 Likert items. Ranking scores range from 1 to 4, with 1 being the highest, or best model, and 4 being the lowest. We can see that rankings and Likert scores averaged across both features increase with the number of training queries used for a model, with a leveling off between 50 and 100 queries. The same trend is observed when individually inspecting the human distance feature. The stove distance feature shows evidence of flatter leveling off, with models trained with 50 and 100 queries having equivalent average Likert scale scores, and the model with 50 queries having a slightly higher average ranking than 100 queries.
Table 4: Mean and sample standard deviation of Likert and Ranking scores for calibrated features trained on different query counts.
Feature Metric 0 Queries 25 Queries 50 Queries 100 Queries
Stove Distance Likert Score 13.67 $\pm$ 6.92 18.25 $\pm$ 5.61 22.58 $\pm$ 3.63 22.58 $\pm$ 4.01
Ranking 3.08 $\pm$ 1.24 3.08 $\pm$ 0.79 1.83 $\pm$ 0.94 2.00 $\pm$ 0.95
Human Distance Likert Score 10.92 $\pm$ 5.57 20.50 $\pm$ 3.45 22.00 $\pm$ 4.65 22.83 $\pm$ 5.54
Ranking 3.50 $\pm$ 1.00 2.75 $\pm$ 0.87 2.17 $\pm$ 1.03 1.58 $\pm$ 0.67
Both Features Likert Score 12.29 $\pm$ 6.30 19.38 $\pm$ 4.70 22.29 $\pm$ 4.09 22.71 $\pm$ 4.73
Ranking 3.29 $\pm$ 1.12 2.92 $\pm$ 0.83 2.0 $\pm$ 0.98 1.79 $\pm$ 0.83
D.2 Preference analysis
Each participant provided a short free-response description of their contextual preference prior to responding to contextual feature queries (see prompt in Fig. 20). Therefore, we collected 12 short free-response answers regarding contextual preferences for each of the two test features. A single reviewer manually reviewed the responses and conservatively grouped them based on the level of specificity provided (Tables 56). Responses describing contextual preferences in general, high-level terms without specific details were grouped together, while responses including concrete details were grouped separately unless the details matched exactly. This inductive grouping was intended to get a preliminary understanding of preference diversity. We found 6 users shared a general preference for the human distance feature, with 2 users sharing a similar variant and 4 users holding unique variants. For the stove distance feature, we found 7 users shared a general preference, and the remaining 5 users described individual variations.
Figure 17: Differences in user preferences captured in learned calibrated features.
In Fig 17 we give another example similar to Fig 6, where we demonstrate how learned calibrated features differ for users with different preferences. The main difference in this example is the context in which the stove is off – while for one user the feature is equally respected in all states when the stove is off, the other user still desires the robot to stay away from directly above the stove, even when it is off.
D.3 Environments
The same environments are used, specifically the Cup environment was used for the training feature and the Utensil environment was used for the two test features, however with discrete contexts for visualization purposes as defined at the end of App.B. Environment state sets of 10k states (with discrete contexts) were generated, and a set of 100 queries (50 for training feature) were randomly sampled and maintained consistent across participants. The same state pairs were shown for training both test features in the utensil environment. Additionally, while the laptop location was represented in the state, when training the two test features in the Utensil environment, the irrelevant laptop was not visualized in order to prevent cluttering of the visualization.
Figure 18: Query response GUI. State 1 shown for first state pair in comparison set, which was the same across both experiment features. Participants visualized and responded to queries using buttons on right panel. Left hand panel is irrelevant and can be ignored.
Figure 19: Model visualization GUI. Example shown of calibrated feature trained for stove distance with 50 contextual feature queries. Visualized in high heat context. Participants inspected different models using buttons on right panel. Left hand panel is irrelevant and can be ignored.
D.4 Graphical user interfaces
Participants interacted with two graphical user interfaces (GUIs) in order to both respond to contextual feature queries (Fig. 18), and to interactively inspect learned calibrated features (Fig. 19). When responding to queries, participants were able to view the states from different angles, and in response to the visualized contextual feature query select between "Select State 1", "Select State 2", or "States are equal". Users could also switch back and forth between the two compared states, and the current state was visualized in order to help users stay aware. Additionally, environment objects, like utensils, were colored differently to help make relevant context immediately salient and speed up query response.
When visualizing and interactively inspecting calibrated features, participants were presented with four models ("Relationships" 1-4) which represented calibrated features trained on 0 (base feature), 25, 50, and all 100 labeled queries. Learned models were visualized just as described in App. C.5, however 5000 states were used as opposed to 8000 for speed of evaluation and visualization. Participants could switch between the four models, and visualize a model in varying contexts by dragging a sliding bar which snapped the visualized context to the nearest of four spread out discrete contexts (and changed the model visualization accordingly). Participants were instructed on helpful ways to investigate and compare models, such as inspecting a single models across contexts, and switching between all models to see how they compare in a single context. Participants provided no input in this GUI, instead using it to respond to a survey collecting the Likert scale evaluation for each model and model rankings.
D.5 Procedures
Participants completed a familiarization phase, where they were trained in contextual feature queries and practiced providing 50 contextual feature queries for the laptop distance feature in the Cup environment. Participants also practiced visualizing and inspecting three learned models trained on 0 (base feature), 25, and all 50 of their provided practice queries. Participants then completed two experiment rounds of query response, which only differed from the practice round in the number of queries answered (100), and the resulting number of models visualized (4 instead of 3).
At the start of each query response round, participants were primed with a general preference, prompting them to focus on a specific contextual element to teach the effect of. Participants were then asked to write down their personal contextual preference (see prompt example for stove distance feature in Fig. 20). Again, while our method supports accounting for the contextual effects of numerous contextual elements, we used this prompting to steer participants to teach the contextual effects of a single element such that we could visualize changes in the affecting contextual element when they visualized the learned models. The query response GUI was then pulled up (Fig. 18), and participants were prompted with a contextual feature query (Fig.  21) which remained constant during the query round.
After completing a round of query response, participants completed post-query questionnaires and inspected the learned models (Fig. 19. Three calibrated feature models were trained while the participant took a 1-2 minute break. The participant then inspected the three learned models and the base feature, responded to the Likert scale for each model, and ranked the four models.
The full experiment script, training presentation, and survey can be found in supplementary material. We provide one survey ordering, where in the other the experiment features are swapped.
Figure 20: Example contextual preference prompt. Short answer descriptions of personal preferences were collected. Despite general priming, participants showed variety in responses.
Figure 21: Example contextual feature query language. Query stayed constant throughout round and was displayed next to query response GUI.
Feature order counterbalancing. We counterbalanced the ordering of the two experimental features: half of the participants saw one feature first, and the other half saw the other feature first. When changing the order of these features, we kept the mapping between the calibrated feature values (trained at 0, 25, 50, and 100 for non-training) and their corresponding numbered relationships (1, 2, 3, and 4 for non-training) fixed. There were 6 possible orderings of the first training feature. For each of these, there were 24 possible combinations of the second and third training features—resulting in too many total permutations to fully counterbalance across all three. To manage this, we used all 6 possible orderings of the first training feature, with each ordering appearing twice (once in each of the two experimental feature orderings). The orderings of the second and third training features were selected to be unique across participants within each ordering condition, but were reused across the two conditions. We aimed for approximate balance by rotating which training queries appeared in each position and ensuring that each participant saw a diverse set of training query orders.
Table 5: Grouped preferences for human distance feature.
Group ID Group Description Free Response Preference
1 General preference (n=6) "I would like to make sure that a sharper utensil is further from me as possible. A dull utensil is ok if it is closer."
1 General preference (n=6) "The sharper the utensil, the further it should be from the person."
1 General preference (n=6) "I would like sharper objects to be further away from the person and less sharp objects can be closer."
1 General preference (n=6) "The sharper the utensil the further away"
1 General preference (n=6) "I think I would like sharper utensils to be farther away from the human."
1 General preference (n=6) "Utensil sharpness should correlate with the preferred distance to the person."
2 Unique preference "Keep the knife very far away from human, fork can be at half the arms length. Spatula and spoon could be close but not too close (a hands length away)"
3 Unique preference "The spoon and spatula should be fine wherever. I’d like the fork and knife to be outside a range around the person, after which I don’t really care where they are."
4 Unique preference "There should be a general safety zone around the near vicinity of the person that no utensil should cross (personal space). The sharper the object, the larger that safety zone should be."
5 Unique preference "Distance for the spoon and spatula doesn’t matter. The knife needs to be kept further than the fork but past a certain distance it doesn’t matter how far they are."
6 Specific shared preference (n=2) "generally, the sharper the utensil the further away from the person it should be. Spoons and spatulas are roughly equal in sharpness."
6 Specific shared preference (n=2) "Sharper objects should be further away from the human. The spoon and spatula are equally sharp"
Table 6: Grouped preferences for stove distance feature.
Group ID Group Description Free Response Preference
1 General preference (n=7) "the robot’s hand should be father away from the stove as the heat increases."
1 General preference (n=7) "If the stove is hot, I would like the utensil to be further away. If the stove is not hot, the utensil can be closer."
1 General preference (n=7) "The hotter the stove is the further away the hand should be"
1 General preference (n=7) "The robot hand should be further away from a hotter stove"
1 General preference (n=7) "Increasing stove heat should correlate with the preferred distance from the end effector to the stove."
1 General preference (n=7) "If the stove is on, the utensil should not be directly over it; if the stove is off, this is OK. If the stove is hot, the utensil should not be in relative proximity (tapers off with temperature)."
1 General preference (n=7) "I would want the robot hand to be further away from the stove the hotter the stove is."
2 Unique preference "If the stove heat is at the highest keep the EE furthest away. If the stove is on in general It should be a hands length away. If it’s off the EE can be anywhere"
3 Unique preference "If the stove is on at all, the robot should be at least half the countertop width away."
4 Unique preference "Disregarding what material the utensil is made out of, I would say that the utensil (context) does not really matter. There is a circle around the stove, inside of which I would not like the robot hand. Outside of this circle, I don’t really care where the robot hand is."
5 Unique preference "I don’t want the robot to be close to the stove. Past a certain distance, I don’t care where the robot is. Close is probably where I would expect the robot to feel the heat."
6 Unique preference "I would like the robot’s hand to be further away from the stove the hotter the stove is except when there’s a spatula in hand."